Antuan Company set the following standard costs for one unit of its product.

Direct materials (4.0 Ibs. @ $6.00 per Ib.) $ 24.00
Direct labor (1.8 hrs. @ $12.00 per hr.) 21.60
Overhead (1.8 hrs. @ $18.50 per hr.) 33.30
Total standard cost $ 78.90


The predetermined overhead rate ($18.50 per direct labor hour) is based on an expected volume of 75% of the factory’s capacity of 20,000 units per month. Following are the company’s budgeted overhead costs per month at the 75% capacity level.

Overhead Budget (75% Capacity)
Variable overhead costs
Indirect materials $ 15,000
Indirect labor 75,000
Power

15,000

Repairs and maintenance 30,000
Total variable overhead costs $ 135,000
Fixed overhead costs
Depreciation—Building 24,000
Depreciation—Machinery 71,000
Taxes and insurance 16,000
Supervision 253,500
Total fixed overhead costs 364,500
Total overhead costs $ 499,500


The company incurred the following actual costs when it operated at 75% of capacity in October.

Direct materials (60,500 Ibs. @ $6.10 per lb.) $ 369,050
Direct labor (20,000 hrs. @ $12.20 per hr.) 244,000
Overhead costs
Indirect materials $ 41,050
Indirect labor 176,250
Power 17,250
Repairs and maintenance 34,500
Depreciation—Building 24,000
Depreciation—Machinery 95,850
Taxes and insurance 14,400
Supervision 253,500 656,800
Total costs $ 1,269,850

rev: 03_28_2018_QC_CS-122864

Problem 23-3A Part 1&2

Required:
1&2. Prepare flexible overhead budgets for October showing the amounts of each variable and fixed cost at the 65%, 75%, and 85% capacity levels and classify all items listed in the fixed budget as variable or fixed.
Answer #1
Antuan company
Flexible Overhead budget
Per unit 65 % capacity 75 % capacity 85 % capacity
20000 * 65 % 20000 * 75 % 20000 * 85 %
Units 13000 15000 17000
VOH costs
Indirect material 15000 / 20000 * 75 % = 1 13000 15000 17000
Indirect labor 75000 / 15000 = 5 65000 75000 85000
Power 15000 / 20000 * 75 % = 1 13000 15000 17000
R and M 30000 / 20000 * 75 % = 2 26000 30000 34000
Total VOH costs 117000 135000 153000
FOH costs
Dep. - Building 24000 24000 24000
Dep. - Machinery 71000 71000 71000
Taxes and Ins. 16000 16000 16000
Supervision 253500 253500 253500
Total FOH costs 364500 364500 364500
Total OH costs 481500 499500 517500